how do you solve:
is triangle DEF, d=5 e=6 f=3 then COS E=????
i have the right set up i believe but i cant get the right answer helpp!!!

Answer by venugopalramana(3286)

is triangle DEF, d=5 e=6 f=3 then COS E=????
i have the right set up i believe but i cant get the right answer helpp!!!
FORMULA IS
e^2=d^2+f^2+2df*COS(E)
6^2=5^2+3^2+2*5*3*COS(E)
COS(E)=(36-25-9)/(2*5*3)=1/6